OpenStudy (anonymous):

Auditions for a new, upcoming band in Chicago were held with 124 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 28, how many women auditioned?

OpenStudy (anonymous):

28

OpenStudy (anonymous):

70

OpenStudy (dan815):

hi u have 2 equation

OpenStudy (anonymous):

Lets let x=the number of women. We know that there is a total of 124 people. So, the number of males is 124-x. Now, we know that twice the number of men less three times the number of women is equal to 28, or: \[2(124-x) - 3x=28\] so: \[248-2x-3x=28\] or \[220-5x=0\] so: \[5x = 220\] so x=220/5=44

OpenStudy (anonymous):

or 44

OpenStudy (dan815):